OpenStudy (anonymous):

Solve the system of equations by graphing. 2x-y=-2, 4x+y=-4

OpenStudy (anonymous):

Solve both for y and get it the slope intercept form y=mx+b then graph it.

OpenStudy (anonymous):

that is a hard way to solve a system

OpenStudy (anonymous):

\[2x-y=-2\iff y=2x+2\]\[4x+y=-4\iff y=-4x-4\]

OpenStudy (anonymous):

now set them equal and solve \[2x+2=-4x-4\]